Fermi National Laboratory
 

Accelerator Update for November 26, 2007 - November 28, 2007

 

Monday November 26

The day shift began with Operations monitoring the Tevatron (TeV) store 5750 (100.56E30) and the Recycler stash (62.1E10), with the Antiproton Source (Pbar) stacking (44.2E10), with the Main Injector (MI), Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams, and with the NuMI and SciBooNE / MiniBooNE experiments taking beam.

At 8:17 AM, the NuMI beam permit tripped off due to a power supply trip (E:V118). It reset.

At 8:27 AM, a Linac RF station (LRF3) tripped off for the first of four times in 1.5 hours. It reset.

Booster RF experts brought BRF5 online and put BRF19 in standby at 9:50 AM.

An expert replaced a control card for E:V118 at 1:06 PM.

Operations reported at 3:52 PM that the Recycler had stashed antiprotons three times during the day shift.

MI experts began a coalescing study at 4:32 PM.

Operations terminated store 5750 at 9:58 PM.

Operations put the TeV into shot setup at 10:16 PM.

Operations established store 5751 at 11:56 PM with an initial luminosity of 151.52E30.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Tuesday November 27

The midnight shift began with Operations monitoring store 5751 (134.54E30) and the stash (88.7E10), with Pbar stacking (9.4E10), and with Sci /MiniBooNE taking beam.

At 5:36 AM, CDF reported that the Tevatron Electron Lens (TEL1) was off. It reset.

FESS technicians reported at 7:22 AM that they were preparing to replace the TeV sector B0 pond water pump. Later, they reported that they could not begin this repair.

Operations reported at 8:05 AM that the Recycler had stashed four times during the midnight shift.

Operations reported at 10:13 AM that Pbar set a new stacking record with a one-hour rate of 23.22E10.

At 10:45 AM, Linac experts began replacing a timing circuit for their low level RF system. This job will hold off beam for a while. During this downtime, MI experts began replacing a power supply control card and instrumentation experts removed the notch filter from a MI DCCT and then began a calibration.

Linac experts completed their RF work at 11:15 AM. Operations resumed sending beam to Sci/MiniBooNE.

Pbar resumed stacking and NuMI resumed taking beam at 11:30 AM.

Instrumentation experts completed their MI DCCT calibration at 11:33 AM.

At 2:21 PM, store 5751 aborted. There was no quench. The abort was due to an RF anode power supply at TeV sector F0. Later, experts reported that they had found clock errors for F1 and F2. They continued to track the problem.

Operations prepared D0 for access at 4 PM.

Operations reported at 4:03 PM that the Recycler had stashed once during the day shift.

Controls experts reported at 4:37 PM that they had replaced a receiver and transmitter at TeV sectors F2 and F3.

Around 6 PM, Operations reported that the cryo system at TeV sector C4 was going bad. TeV experts investigated and reported a problem with the wet engine.

D0 came out of access at 6:20 PM.

At 7:15 PM, operators discovered that the C4 wet engine was getting worse. Operators talked with the Cryo Coordinator and turned the engine off. Experts were being called in to begin repairs. Per instruction, operations closed various beam valves at C4 to protect the vacuum for the separators.

Cryo experts arrived at 8:50 PM to begin their work at C4. At 9:56 PM, the Cryo Coordinator reported that all they needed to replace was the flywheel.

At 11:02 PM, the TeV suffered a sector C4 quench. When the C4 wet engine was turned on, it pushed a slug of warm gas through the system and caused the quench.

Cryo experts reported at 11:31 PM that all cryo work had been completed. Operations waited for the cryo systems to recover.

Wednesday November 28

The midnight shift began with Operations monitoring a stash (244.4E10), with Pbar stacking (142.7E10), and with NuMI and Sci/MiniBooNE taking beam.

Operations ramped the TeV at 1:30 AM. At 1:36 AM, the TeV suffered an F11 quench. Operators investigated, but could find no cause for the quench. Operations contacted experts.

Operations reported at 3:45 AM that they still don’t know what caused the quench, but that a number of vacuum problems appeared in TeV sector F1. More experts were called in to help with the investigations.

Operations began preparing the TeV for an access at 5:46 AM to search for vacuum problems. This access would halt beam to Sci/MiniBooNE.

Experts reported that they had found a pulse of beam directed toward the AP0 target, which passed through F11, that didn’t arrive. Investigations continued.

The Plans for Wednesday

The plans for today are to determine what caused this quench and repair any vacuum problems in F sector. The Run Coordinator will allow certain maintenance jobs during these TeV investigations.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/28/2007 -->    email Fermilab